# Dilaton Sum Rules of Gravitational Form Factors in QCD at Order $\alpha_{s}$

### 25 Mar 2026 v1

@@ 1 Introduction
Gravitational form factors (GFFs) provide a compact description of hadron structure in terms of matrix elements of the QCD energy-momentum tensor. They encode momentum fractions, total and orbital angular momentum, and the mechanical properties of hadrons such as pressure and shear forces. Experimentally they are not accessed through dynamical gravity, but through hard exclusive processes, most notably deeply virtual Compton scattering (DVCS), where generalized parton distributions (GPDs) enter the factorized amplitude and connect measurable observables to moments of the stress tensor [1, 2, 3, 4, 5].
The relevant hard subprocesses can be analyzed through correlators involving insertions of the energy-momentum tensor. The basic subleading interaction that we are going to consider is the non-Abelian three-point function $TJJ$, with one stress tensor and two gauge currents, shown as an insertion in Fig.1 (right). This correlator is the natural arena in which the conformal anomaly, its dispersive representation and its interpolation by effective scalar degrees of freedom can be studied perturbatively [6, 7, 8, 9]. The purpose of the present contribution is to summarize this picture and to make explicit the equations underlying the discussion.
@@ 3.1 The reconstruction of the quark and gluon sectors in QCD
The perturbative analysis requires the QCD stress-energy tensor. In the gauge-fixed theory it is obtained by combining the field-strength, fermionic, gauge-fixing and ghost contributions,
@@
where $\mathcal{A}^{\alpha\beta ab}$ denotes the contribution generated by differentiating the anomaly functional with respect to the gauge fields. The first term is fixed by the anomaly functional and is mass independent, whereas the second contains the explicit insertion of the trace of the stress tensor and therefore carries the dependence on quark masses and on the details of the broken phase. This distinction is central when identifying the genuine anomaly form factor away from the strict conformal point, and needs to be taken into account in the analysis of the sum rule. It is conceptually the same separation encountered in the electroweak theory when one isolates the genuine anomalous dilaton–gauge–gauge coupling from ordinary loop-induced scale breaking proportional to the Higgs vacuum expectation value or to explicit fermion masses [17].
@@ 4 Momentum-space CFT and the role of gauge fixing
The $TJJ$ correlator can be organized using the techniques of CFT in momentum space [21, 10, 11, 8, 22, 23]. In a genuine conformal theory, three-point functions of currents and stress tensors are constrained by conformal Ward identities and can be decomposed into transverse-traceless, longitudinal and trace sectors. The elementary projectors are
@@ 7 On-shell limit and pole dominance
which shows explicitly that the anomaly exchange is nonlocal. The character of the interaction, as shown in [14] is to define a light-cone excitation [14]. At the same time, the full $TJJ$ amplitude contains another pole contribution in the traceless sector This second structure allows us to identify an additional contribution to the anomaly effective action describing the interaction around the light-cone, of the form
@@
The $n_{f}$ part was identified in [13] as a possible plasmon mode in QED. In the non-Abelian case the extra contribution is proportional to $(n_{f}-C_{A})$. The situation is therefore slightly different from the $AVV$ or $TTJ_{5}$ cases [26, 27], where the residue of the (single) pole present in the anomaly part is sufficient, using the conformal cosntraints to determine the entire structure of the chiral correlators.
@@ 8 Relation to anomaly effective actions
This observation also clarifies the distinction between the present $TJJ$ analysis and the study of higher stress-tensor correlators such as $TTT$. The latter probe the full gravitational sector of the anomaly and involve a substantially richer set of tensor structures [22, 23]. By contrast, the $TJJ$ correlator isolates the coupling of one stress tensor to two gauge currents and is the natural object for the partonic discussion of GFFs, being the anomaly dependent on the gauge fields. The two problems are related, but they are not equivalent: the first is primarily about hadronic structure and short-distance factorization, whereas the second concerns the full nonlinear realization of the conformal anomaly in a gravitational context, with the inclusion of extraWeyl invariant terms such as $C^{2}$ (the square of the Weyl tensor) and $E$, the Euler-Poincare’ density (Gauss-Bonnet term).
@@ 9 Physical interpretation and relation to hadronic observables
At the same time, one should be precise about the physical meaning of GFFs. They do not measure dynamical gravity. They measure matrix elements of the QCD stress tensor, with the metric treated as an external source. To probe genuine gravitational dynamics one would have to move to higher stress-tensor correlators such as $TTT$, which, as just mentioned, project onto other sectors of the conformal anomaly and lie beyond the immediate scope of the hadronic problem [22, 23]. The present discussion is therefore best viewed as a bridge between two problems: the extraction of hadronic structure from exclusive QCD amplitudes and the field-theoretic interpretation of anomaly-induced exchanges in perturbation theory. In this sense the $TJJ$ correlator occupies a privileged intermediate position. It is simpler than the full purely gravitational sector, but already rich enough to display the emergence of a scalar anomaly channel, its dispersive sum rule and its interplay with the ordinary traceless tensor exchange relevant for partonic scattering.

### 27 Mar 2026 v2

@@ 1 Introduction
Gravitational form factors (GFFs) provide a compact description of hadron structure in terms of matrix elements of the QCD energy-momentum tensor. They encode momentum fractions, total and orbital angular momentum, and the mechanical properties of hadrons such as pressure and shear forces. Experimentally they are not accessed through dynamical gravity, but through hard exclusive processes, most notably deeply virtual Compton scattering (DVCS), where generalized parton distributions (GPDs) enter the factorized amplitude and connect measurable observables to moments of the stress energy tensor [1, 2, 3, 4, 5].
The relevant hard subprocesses can be analyzed through correlators involving insertions of this tensor. The basic subleading interaction that we are going to consider is the non-Abelian three-point function $TJJ$, with one stress energy tensor and two gauge currents, shown as an insertion in Fig.1 (right). This correlator is the natural arena in which the conformal anomaly, its dispersive representation and its interpolation by effective scalar degrees of freedom can be studied perturbatively [6, 7, 8, 9]. The purpose of the present contribution is to summarize this picture and to make explicit the equations underlying the discussion.
@@ 3.1 The reconstruction of the quark and gluon sectors in QCD
The perturbative analysis requires the QCD stress energy tensor. In the gauge-fixed theory it is obtained by combining the field-strength, fermionic, gauge-fixing and ghost contributions,
@@
where $\mathcal{A}^{\alpha\beta ab}$ denotes the contribution generated by differentiating the anomaly functional with respect to the gauge fields. The first term is fixed by the anomaly functional and is mass independent, whereas the second contains the explicit insertion of the trace of the stress energy tensor and therefore carries the dependence on quark masses and on the details of the broken phase. This distinction is central when identifying the genuine anomaly form factor away from the strict conformal point, and needs to be taken into account in the analysis of the sum rule. It is conceptually the same separation encountered in the electroweak theory when one isolates the genuine anomalous dilaton–gauge–gauge coupling from ordinary loop-induced scale breaking proportional to the Higgs vacuum expectation value or to explicit fermion masses [17].
@@ 4 Momentum-space CFT and the role of gauge fixing
The $TJJ$ correlator can be organized using the techniques of CFT in momentum space [21, 10, 11, 8, 22, 23]. In a genuine conformal theory, three-point functions of currents and stress energy tensors are constrained by conformal Ward identities and can be decomposed into transverse-traceless, longitudinal and trace sectors. The elementary projectors are
@@ 7 On-shell limit and pole dominance
which shows explicitly that the anomaly exchange is nonlocal. The character of the interaction, as shown in [14] is to define a light-cone excitation [14]. At the same time, the full $TJJ$ amplitude contains another pole contribution in the traceless sector. This second structure allows us to identify an additional contribution to the anomaly effective action describing the interaction around the light-cone, of the form
@@
The $n_{f}$ part was identified in [13] as a possible plasmon mode in QED. In the non-Abelian case the extra contribution is proportional to $(n_{f}-C_{A})$. The situation is therefore slightly different from the $AVV$ or $TTJ_{5}$ cases [26, 27], where the residue of the (single) pole present in the anomaly part is sufficient, using the conformal cosntraints, to determine the entire structure of the chiral correlators.
@@ 8 Relation to anomaly effective actions
This observation also clarifies the distinction between the present $TJJ$ analysis and the study of higher stress energy tensor correlators such as $TTT$. The latter probe the full gravitational sector of the anomaly and involve a substantially richer set of tensor structures [22, 23]. By contrast, the $TJJ$ correlator isolates the coupling of one stress energy tensor to two gauge currents and is the natural object for the partonic discussion of GFFs, being the anomaly dependent on the gauge fields. The two problems are certainly related, but they are conceptually distinct. The first is mainly concerned with the gauge structure, while the second involves the full nonlinear realization of the conformal anomaly in a gravitational framework, with the inclusion of additional Weyl-invariant terms such as $C^{2}$, the square of the Weyl tensor, and $E$, the Euler–Poincaré density, or Gauss–Bonnet term, absent in the GFF case.
@@ 9 Physical interpretation and relation to hadronic observables
At the same time, one should be precise about the physical meaning of GFFs. They do not measure dynamical gravity. They measure matrix elements of the QCD stress energy tensor, with the metric treated as an external source. To probe genuine gravitational dynamics one would have to move to higher stress-tensor correlators such as $TTT$, which, as just mentioned, project onto other sectors of the conformal anomaly and lie beyond the immediate scope of the hadronic problem [22, 23]. The present discussion is therefore best viewed as a bridge between two problems: the extraction of hadronic structure from exclusive QCD amplitudes and the field-theoretic interpretation of anomaly-induced exchanges in perturbation theory. In this sense the $TJJ$ correlator occupies a privileged intermediate position. It is simpler than the full purely gravitational sector, but already rich enough to display the emergence of a scalar anomaly channel, its dispersive sum rule and its interplay with the ordinary traceless tensor exchange relevant for partonic scattering.